# Changing the ranking in eigenvector centrality of a weighted graph by small perturbations

### v2

@@ 3. Robustness and vulnerability of eigenvector centrality
Given a graph $\widetilde{\mathcal{G}}=(\mathcal{V},\mathcal{E},\widetilde{A})$, with the same vertices and edges of $\mathcal{G}$, we define the distance between $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ by
@@
$A=\left(\begin{array}[]{ccccc}0&14.9&6.7&3.9\\ 0&0&4.0&0\\ 0&5.2&0&14.1\\ 4.8&0&9.6&0\end{array}\right),\quad\mathrm{v}=\left(\begin{array}[]{c}0.5665\\ 0.1570\\ \mathbf{0.5844}\\ 0.5594\end{array}\right).\hskip 2.84544pt$
@@ 3.1. Robustness of the ranking and nearest ambiguous graph
We consider the general case of a directed graph and look for the closest strongly connected graph $\widetilde{\mathcal{G}}=(\mathcal{V},\mathcal{E},\widetilde{A})$ to $\mathcal{G}=(\mathcal{V},\mathcal{E},A)$, such that the $m$ highest entries of the Perron eigenvector coalesce. In order to preserve the strong connectivity we have to impose that the perturbed graph preserves the property $\widetilde{a}_{ij}>0$ for $(i,j)\in\mathcal{E}$. For computational purposes we replace it with the property
@@
with $01\gg\delta>$ 0 a small positive tolerance.
@@ 4.1. First order perturbation theory for simple eigenvalues
The following standard perturbation result for eigenvalues is well-known; see e.g. [14, Theorem 1] and [23, Section II.1.1].
@@
Consider a continuously differentiable path of square real matrices $M(t)$ for $t$ in an open interval $I$. Let $\lambda(t)$, $t\in I$, be a continuous path of real simple eigenvalues of $M(t)$. Let $x(t)$ and $y(t)$ be left and right eigenvectors, respectively, of $M(t)$ corresponding to the eigenvalue $\lambda(t)$. Then, $x(t)^{\top}y(t)\neq 0$ for $t\in I$ and $\lambda$ is continuously differentiable on $I$ with the derivative (we denote ˙$\dot{\phantom{a}}=d/dt$)
@@
(10) $\|x(t)\|=1,\ \ \|y(t)\|=1,\quad x(t)^{\top}y(t)\text{ is real and positive.}$.
@@ 4.2. First order perturbation theory for eigenvectors
For the derivative of eigenvectors we need the notion of group inverse (or reduced resolvent); see [27] as well as Kato ([23, Section I.5.3]).
@@
It is known from Meyer and Stewart ([27]) that if $M$ is normal, then its group inverse $M^{\#}$ is equal to the better known Moore–Penrose pseudoinverse $M^{\dagger}$. In general, the two pseudoinverses are not the same. They are, however, related by the following result, which is a special case of a more general result in [18, Appendix A] which is also verified directly.
@@
Suppose that the real matrix $M$ has the simple eigenvalue 00 with corresponding left and right real eigenvectors $x$ and $y$ of unit norm and such that $x^{\top}y>0$. Let $M^{\#}$ be the group inverse of $M$, and with $\kappa=1/(x^{\top}y)$ define the projection $\Pi={\rm I}-\kappa xy^{\top}$. The group inverse $M^{\#}$ of $M$ is related to the Moore–Penrose pseudoinverse $M^{\dagger}$ by
@@ 5. A methodology for computing the closest ambiguous graph and the robustness radius
[…]
@@ 5.3. Constrained gradient of the functional
To determine the admissible direction $\dot{E}$ of steepest descent from $E$, we therefore consider the following optimization problem for $G=G_{\varepsilon}(E)$:
@@
(20) $\begin{cases}\langle E,Z\rangle=0\,,\\ z_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,\\ z_{ij}=z_{ji}\text{ for all }(i,j)\quad\text{(for undirected graphs})\,,\\ \langle Z,Z\rangle=1\,.\end{cases}$
@@
(22) $\begin{cases}\langle E,\dot{E}\rangle=0\,,\\ \dot{E}_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,\\ z_{ij}=z_{ji}\text{ for all }(i,j)\quad\text{(for undirected graphs})\,,\\ \langle G,Z\rangle=-1\,.\end{cases}$
@@
$\kappa Z=-G-\gamma E+\sum_{(i,j)\in\mathcal{E}_{0}}\mu_{ij}{\rm u}_{i}{\rm u}_{j}^{T}\,,$
$Z_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,$
$\langle E,Z\rangle=0\,,$
$\langle Z,Z\rangle=1\,,$
@@
$\mu_{ij}z_{ij}=0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,$
$\mu_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,.$
where ${\rm u}_{\ell}$ is the $\ell$-th unit versor, $\gamma$ is determined by the normalization $\langle E,Z\rangle=0$ and $\kappa>0$ is determined by the normalization $\lVert Z\rVert^{2}=\langle Z,Z\rangle=1$.
@@ 5.6. Underlying low-rank property of extremizers
where $G=G_{\varepsilon}(E)$. In this case extremizers (stationary points) are characterized by an underlying low-rank property, being $G$ the projection of a low-rank matrix onto the sparsity-pattern of the weighted adjacency matrix. Clearly, if $\varepsilon<\min_{ij}a_{ij}$, inequality constraints cannot be active so that we can ignore them and solve the following system of ODEs (28). Moreover, in all our experiments we have never observed activation of inequality constraints associated to nonnegativity.
@@ 6.1. An illustrative example of an undirected graph
Modifying the full pattern of $A$
@@
Modifying only a subpattern of $A$
As already mentioned in the discussion following Definition 5.1, it is possible to consider, instead of the whole set of edges, a subset ${\mathcal{E}}^{\prime}\subset\mathcal{E}$ and proceed exactly as discussed by simply replacing the projection $P_{\mathcal{E}}$ with $P_{\mathcal{E}^{\prime}}$. As an example we consider again the simple undirected graph with $9$ vertices of Figure 1 and modify only the edges not incident to the two nodes $v_{1}$ and $v_{4}$. The results so obtained are illustrated in Figure 3. We compute the Perron eigenvector $\mathrm{v}$
@@ 6.2. Numerical integration of the gradient system
then we reduce the step size for the next step to $h/\theta$. If the step size has not been reduced in the previous step, we then increase the step-length to be used in $[t_{k+1},t_{k+2}]$.
In a numerical solution of the gradient system (for example using the Euler-based scheme (34)), we have to monitor the sets of edges where $a_{ij}+\varepsilon e_{ij}=\delta$ and among those we must further track those edges where the sign of $-g_{ij}+\gamma e_{ij}$ changes (where $g_{ij}$ is the $(i,j)$-th entry of the matrix $G_{\varepsilon}(E)$ and $e_{ij}$ is the $(i,j)$-th entry of the matrix $E$). When the active set $\{(i,j)\,:\,a_{ij}+\varepsilon e_{ij}=\delta\}$ is changed, then $\gamma$ also changes.
@@ 6.3. Outer iteration
Let $E(\varepsilon)$ denote the minimizer of the functional $F_{\varepsilon}$. Generically we expect that for a given perturbation size $\varepsilon<\widetilde{\varepsilon}_{m}$, $E(\varepsilon)$ is smooth, which means that the spectral radius of $A+\varepsilon E$ is a simple eigenvalue. If so, then $f(\varepsilon):=F_{\varepsilon}(E(\varepsilon))$ is a smooth function of $\varepsilon$ and we can exploit its regularity to obtain a fast iterative method to converge to $\widetilde{\varepsilon}_{m}$ from the left. Otherwise we can use a bisection technique to approach $\widetilde{\varepsilon}_{m}$.
An example is provided in Figure 4.
@@
(38) $\hskip-25.60747ptf^{\prime}(\varepsilon)=\Bigl{\langle}G_{\varepsilon}(E(\varepsilon)),E(\varepsilon)+\varepsilon E^{\prime}(\varepsilon)\Bigr{\rangle}=\Bigl{\langle}G_{\varepsilon}(E(\varepsilon)),E(\varepsilon)\Bigr{\rangle}.$
@@ 7.0.1. Example 1, matrix $A_{27}$
$k$ $\varepsilon_{k}$ $f_{c}(\varepsilon_{k})$#\## eigs
@@
[…]
@@
[…]
@@
In a first experiment we have applied the Matlab routine fmincon to the problem formulation (4)–(5) for the case $m=2$. Here we use fmincon as a black box, without providing the analytic form of the gradient (the derivation of which is among the results of this paper) to the routine. Using standard default values we have not been able to obtain a correct solution, hence it has been necessary to increase the number of function evaluations allowed. Table 2 shows the results of the routine with an interior point algorithm, which is the one that performs best. The resulting solution is very close to the one we found, with coalescent entries #$\#111$ and #$\#315$,
@@
In Table 2, $k$ stands for the iteration number, $x$ is the vector containing the nonzero entries of the upper triangular part of the weighted adiacency matrix $A$, $g(x)$ is the constraint, $\mathrm{v}_{1}-\mathrm{v}_{2}$ being $\mathrm{v}(x)$ the Perron eigenvector of the matrix $B=A-X$, where $X$ is the perturbation associated to $x$. Finally, $f(x)=\sqrt{2}\|x\|=\|X\|_{F}$ and #$\#fe$ indicates the number of function evaluations. The overall CPU time is about $80$ minutes.
$k$#\## fe $f(x_{k})$ $g(x_{k})$
@@
7.0.2. Example $2$, matrix nopoly

### v3

@@ 3. Robustness and vulnerability of eigenvector centrality
[…]
@@
$A=\left(\begin{array}[]{ccccc}0&14.9&6.7&3.9\\ 0&0&4.0&0\\ 0&5.2&0&14.1\\ 4.8&0&9.6&0\end{array}\right),\quad\mathrm{v}=\left(\begin{array}[]{c}0.5665\\ 0.1570\\ \mathbf{0.5844}\\ 0.5594\end{array}\right).\;\$
@@ 3.1. Robustness of the ranking and nearest ambiguous graph
[…]
@@
with $1\gg\delta>0$ a small positive tolerance.
@@ 4.1. First order perturbation theory for simple eigenvalues
The following standard perturbation result for eigenvalues is well-known; see e.g. [14, Theorem 1] and [25, Section II.1.1].
@@
Consider a continuously differentiable path of square real matrices $M(t)$ for $t$ in an open interval $I$. Let $\lambda(t)$, $t\in I$, be a continuous path of real simple eigenvalues of $M(t)$. Let $x(t)$ and $y(t)$ be left and right eigenvectors, respectively, of $M(t)$ corresponding to the eigenvalue $\lambda(t)$. Then, $x(t)^{\top}y(t)\neq 0$ for $t\in I$ and $\lambda$ is continuously differentiable on $I$ with the derivative (we denote $\dot{\phantom{a}}=d/dt$)
@@
(10) $\|x(t)\|=1,\ \ \|y(t)\|=1,\quad x(t)^{\top}y(t)\text{ is real and positive.}$
@@ 4.2. First order perturbation theory for eigenvectors
For the derivative of eigenvectors we need the notion of group inverse (or reduced resolvent); see [29] as well as Kato ([25, Section I.5.3]).
@@
It is known from Meyer and Stewart ([29]) that if $M$ is normal, then its group inverse $M^{\#}$ is equal to the better known Moore–Penrose pseudoinverse $M^{\dagger}$. In general, the two pseudoinverses are not the same. They are, however, related by the following result, which is a special case of a more general result in [19, Appendix A] which is also verified directly.
@@
Suppose that the real matrix $M$ has the simple eigenvalue 0 with corresponding left and right real eigenvectors $x$ and $y$ of unit norm and such that $x^{\top}y>0$. Let $M^{\#}$ be the group inverse of $M$, and with $\kappa=1/(x^{\top}y)$ define the projection $\Pi={\rm I}-\kappa xy^{\top}$. The group inverse $M^{\#}$ of $M$ is related to the Moore–Penrose pseudoinverse $M^{\dagger}$ by
@@ 5. A methodology for computing the closest ambiguous graph and the robustness radius
the Frobenius inner product, i.e. the inner product inducing the Frobenius norm $\|\cdot\|=\|\cdot\|$ on $\mathbb{R}^{n\times n}$.
@@ 5.3. Constrained gradient of the functional
In order to determine the admissible direction $Z=\dot{E}$ of steepest descent from $E$, we therefore consider the following optimization problem for $G=G_{\varepsilon}(E)$:
@@
[…]
@@
[…]
@@
where $\rm u$$_{\ell}$ is the $\ell$-th unit versor, $\gamma$ is determined by the normalization $\langle E,Z\rangle=0$, and $\kappa>0$ is determined by the normalization $\lVert Z\rVert^{2}=\langle Z,Z\rangle=1$.
@@ 5.6. Underlying low-rank property of extremizers
where $G=G_{\varepsilon}(E)$. In this case extremizers (stationary points) are characterized by an underlying low-rank property, being $G$ the projection of a low-rank matrix onto the sparsity-pattern of the weighted adjacency matrix. Clearly, if $\varepsilon<\min_{ij}a_{ij}$, inequality constraints cannot be active so that we can ignore them and solve the system of ODEs (28). Moreover, in all our experiments we have never observed activation of inequality constraints associated to nonnegativity.
@@ 6.1. An illustrative example of an undirected graph
Modifying the full pattern of AA
@@
Modifying only a subpattern of AA
As already mentioned in the discussion preceding Definition 5.1, it is possible to consider, instead of the whole set of edges, a subset $\mathcal{E}$$^{\prime}\subset\mathcal{E}$ and proceed exactly as discussed by simply replacing the projection $P_{\mathcal{E}}$ with $P_{\mathcal{E}^{\prime}}$. As an example we consider again the simple undirected graph with $9$ vertices of Figure 1 and modify only the edges not incident to the two nodes $v_{1}$ and $v_{4}$. The results so obtained are illustrated in Figure 3. We compute the Perron eigenvector $\mathrm{v}$
@@ 6.2. Numerical integration of the gradient system
then we reduce the step size for the next step to $h/\theta$. If the step size has not been reduced in the previous step, we then increase the step-length to be used in $t_{k+1},t_{k+2}]$.
In a numerical solution of the gradient system (for example using the Euler-based scheme (34)), we have to monitor the sets of edges where $a_{ij}+\varepsilon e_{ij}=\delta$ and among those we must further track those edges where the sign of $g_{ij}+\gamma e_{ij}$ changes (where $g_{ij}$ is the $(i,j)$-th entry of the matrix $G_{\varepsilon}(E)$ and $e_{ij}$ is the $(i,j)$-th entry of the matrix $E$). When the active set $\{(i,j)\,:\,a_{ij}+\varepsilon e_{ij}=\delta\}$ is changed, then $\gamma$ also changes.
@@ 6.3. Outer iteration
Let $E(\varepsilon)$ denote the minimizer of the functional $F_{\varepsilon}$. Generically we expect that for a given perturbation size $\varepsilon<\widetilde{\varepsilon}_{m}$, $E(\varepsilon)$ is smooth, which means that the spectral radius of $A+\varepsilon E$ is a simple eigenvalue. If so, then $f(\varepsilon):=F_{\varepsilon}(E(\varepsilon))$ is a smooth function of $\varepsilon$ and we can exploit its regularity to obtain a fast iterative method to converge to $\widetilde{\varepsilon}_{m}$ from the left. Otherwise we can use a bisection technique to approach $\widetilde{\varepsilon}_{m}$. An example is provided in Figure 4.
@@
(38) $\thinspace f^{\prime}(\varepsilon)=\Bigl{\langle}G_{\varepsilon}(E(\varepsilon)),E(\varepsilon)+\varepsilon E^{\prime}(\varepsilon)\Bigr{\rangle}=\Bigl{\langle}G_{\varepsilon}(E(\varepsilon)),E(\varepsilon)\Bigr{\rangle}.$
@@ 7.1. Example 1, matrix $A_{27}$
$k$ $\varepsilon_{k}$ $f_{c}(\varepsilon_{k})$ #\# eigs
@@
We also consider coalescence of the first $m=3$ entries, for which we obtain a perturbation of norm $\widetilde{\varepsilon}_{3}=0.010718306874804$ with the $3$ largest entries of Perron eigenvector $\mathrm{v}$ given by
@@
To obtain $4$ digits of precision we required a smaller tolerance to step the outer iteration, which determined a CPU time of about $9.33$ seconds. Finally we set $m=4$ for which we obtain $\widetilde{\varepsilon}_{4}=0.014389966738210$, with the largest entries of the Perron eigenvector $\mathrm{v}$ being
@@
In a first experiment we have applied the Matlab routine fmincon to the problem formulation (4)–(5) for the case $m=2$. Here we use fmincon as a black box, without providing the analytic form of the gradient (the derivation of which is among the results of this paper) to the routine. Using standard default values we have not been able to obtain a correct solution, hence it has been necessary to increase the number of function evaluations allowed. Table 2 shows the results of the routine with an interior point algorithm, which is the one that performs best. The resulting solution is very close to the one we found, with coalescent entries #111\#111 and #315\#315,
@@
In Table 2, $k$ stands for the iteration number, $x$ is the vector containing the nonzero entries of the upper triangular part of the weighted adiacency matrix $A$, $g(x)$ is the constraint, $\mathrm{v}_{1}-\mathrm{v}_{2}$ being $\mathrm{v}(x)$ the Perron eigenvector of the matrix B=AXB=A-X, where $X$ is the perturbation associated to $x$. Finally, $f(x)=\sqrt{2}\|x\|=\|X\|_{F}$ and #fe\#fe indicates the number of function evaluations. The overall CPU time is about $80$ minutes.
$k$ #\# fe $f(x_{k})$ $g(x_{k})$
@@
7.2. Example $2$, matrix nopoly